A small market orders copies of a certain magazine for its magazine rack each week. Let X =...

x 1 2 3 4 5 6

p(x) 2/16 1/16 3/16 5/16 3/16 2/16


Suppose the store owner actually pays $2.00 for each copy of the magazine and the price to customers is $4.00. If magazines left at the end of the week have no salvagevalue, is it better to order three or four copies of the magazine? [Hint: For both three and four copies ordered, express net revenue as a function of demand X, andthen compute the expected revenue.]

What is the expected profit if three magazines are ordered? (Round your answer to two decimal places.)
$________
What is the expected profit if four magazines are ordered? (Round your answer to two decimal places.)
$________
Answer #1

If 3 magazines are ordered
Total expected revenue = (2/16*1 +1/16*2 + 3/16*3)*4 =$3.25
Cost incurred = $3*2 = $6
Expected profit = $3.25-6 = -$2.75

If 4 magazines are ordered
Total expected revenue = (2/16*1 +1/16*2 + 3/16*3 + 5/16*4)*4 =$8.25
Cost incurred = $4*2 = $8
Expected profit = $8.25-8 = $0.25